OpenStudy (rmrjr22):

determine by inspection at least two solutions of the given first-order IV y'= 3y^(2/3) y(0) = 0

ganeshie8 (ganeshie8):

looks like a trick question

ganeshie8 (ganeshie8):

if initial values are given, then there can exist oly one solution right ?

OpenStudy (rmrjr22):

there are two... so the book says they ended up making it dy/3y^(2/3) = dx

ganeshie8 (ganeshie8):

thats okay, but i dont see yet how two solutions can be there uhmm

ganeshie8 (ganeshie8):

lets dive in and see wat we get maybe..

ganeshie8 (ganeshie8):

next, integrate both sides

ganeshie8 (ganeshie8):

\(\large y'= 3y^{\frac{2}{3}} \)

ganeshie8 (ganeshie8):

its same as : \(\large \frac{dy}{dx}= 3y^{\frac{2}{3}} \)

ganeshie8 (ganeshie8):

cross multiply

ganeshie8 (ganeshie8):

\(\large \frac{dy}{3y^{\frac{2}{3}} }= dx \)

ganeshie8 (ganeshie8):

integrate both side

ganeshie8 (ganeshie8):

\(\large \int \frac{dy}{3y^{\frac{2}{3}} }= \int dx \)

ganeshie8 (ganeshie8):

\(\large \frac{1}{3}\int y^{-\frac{2}{3}} dy= \int dx \)

ganeshie8 (ganeshie8):

\(\large \frac{1}{3} \frac{y^{-\frac{2}{3}+1}}{-\frac{2}{3}+1}= x + c\)

ganeshie8 (ganeshie8):

simplify a bit

ganeshie8 (ganeshie8):

\(\large y^{\frac{1}{3}}= x + c\)

OpenStudy (rmrjr22):

thats it?

ganeshie8 (ganeshie8):

if u plugin (0, 0), u wud get c = 0 so the solution wud be : \(\large y^{\frac{1}{3}}= x \) \(\large y= x^3 \)

ganeshie8 (ganeshie8):

thats the oly one final solution

ganeshie8 (ganeshie8):

you're talking about TWO solutions ?

ganeshie8 (ganeshie8):

*ur textbook

OpenStudy (rmrjr22):

yeah it says at least two... so idk

ganeshie8 (ganeshie8):

uhmm u dont have answers ?

OpenStudy (rmrjr22):

i have y = 0 and y = x^3